# A Comprehensive Survey on Knowledge Distillation

### v1

@@ V-D Speech Input
Feature-based Wang et al.[431], Liu et al.[432], V2I-DETR [433], MaskAgain[434]
Similarity-based RSKD[435], Bhardwaj et al.[436], OOKD [437], DL-DKD[438]
Logit-based Camarena et al.[439], Wu et al.[440], PKD [441], MobileVOS[442], V2I-DETR[433], DTO[443], MVD [444], RankDVQA-mini[445], VideoAdviser [446], Afouras et al.[447], Perez et al.[448]
V-E Video Input
Knowledge distillation has been widely applied in image and language tasks, but its potential in video-based applications is gaining increasing attention. Recent studies have explored KD for video action recognition [432, 440, 439, 431, 441], video classification [436, 435], video object segmentation [442, 433], video instance segmentation [437], Partially Relevant Video Retrieval (PRVR) [438], trajectory forecasting [443], and representation learning through masked video modeling [434, 444]. Table VIII summarizes video distillation methods.
The main challenge in deep learning models, particularly in applications such as video quality estimation [445] and attention prediction [449], is the high computational cost and complexity of large models, which restrict their execution on edge devices. KD effectively addresses this by enabling a smaller, more efficient model (student) to learn from a larger, complex model (teacher) with minimal performance reduction while significantly decreasing computational and memory requirements. MobileVOS [442] specifically focuses on semi-supervised video object segmentation on resource-constrained devices, for example mobile phones, and proposes a pixelwise representation distillation loss to efficiently transfer structural information from the teacher to the student while ensuring feature consistency across frames. In [437], the authors propose a real-time approach that distills similarity information from an offline teacher model, which processes entire video sequences, to an online model, which processes individual frames. [450] introduces an online approach that applies feature distillation on live videos for low-cost semantic segmentation.
Furthermore, KD is applicable in multi-modal learning, enabling cross-modal knowledge transfer, such as video-to-image [433], video-to-text [446], text-to-video [438], audio-to-video [447], and video-to-audio [448]. Notably, [451] leverages KD in a spatio-temporal graph model for the video captioning task, making the model more robust against spurious correlations. V2I-DETR enhances medical video lesion detection by distilling temporal knowledge from a video-based teacher model to an image-based student model, achieving real-time inference speed with high accuracy [433].
VI Applications
@@
VI-A Distillation in Large Language Models
[…]
@@ VI-A Distillation in Large Language Models
LLMs have recently revolutionized the field of NLP, with significant applications across various domains in both academia and industry. However, these models pose critical challenges due to their large number of parameters, which limits their deployment in real-time scenarios. Additionally, these giant models, with a considerable number of attention blocks, have been shown to be overparameterized [3, 4]. Therefore, KD plays a crucial role in compressing LLMs into Small Language Models (SLMs). Existing KD methods for LLMs are mainly classified into two major categories: white-box distillation and black-box distillation. In white-box KD, logits and intermediate outputs of the LLMs are accessible, enabling the student to benefit from the internal information of the teacher. This information can be either logits [10] or intermediate features [22], such as outputs of each attention block or attention scores. On the other hand, in black-box distillation, only the predictions of the teacher are available for distillation, which is common in many closed-source LLMs like GPT-4 [487] and Gemini [488], limiting the distillation process to only the predictions. Black-box methods are categorized into In-Context Learning (ICL) [484], Chain of Thought (CoT) [466], and Instruction Following (IF) [478]. Figure 8 summarizes the black-box and white-box distillation methods for LLMs.
@@
Most white-box KD methods focus on the BERT [2] model [455, 452, 453, 454, 456, 457, 458, 459]. DistilBERT [453] was one of the works that initialized the student with a pre-trained teacher and minimized the soft probabilities between the student and the teacher during the pre-training phase. In a concurrent work, MiniLM [457] proposed distilling the self-attention module of the last transformer layer of the teacher. MobileBERT [456] first trains a special teacher and then uses feature and attention transfer to distill knowledge to the student. TinyBERT [452] introduced a KD method for both pre-training and task-specific training of a smaller version of BERT by distilling the logits, attention matrices, hidden states, and the output of the teacher’s embedding layer. PKD [454] introduces two strategies for incremental distillation, while [458] and [459] propose architecture-agnostic and task-agnostic distillation methods, respectively.
@@
Recently, several new white-box KD methods have been proposed [460, 462, 461, 464, 463, 465], primarily designed for decoder-based LLMs. MetaDistill [460] utilizes a feedback mechanism within a meta-learning framework, GKD [462] trains the student on its self-generated output sequences, and PC-LoRA [464] concurrently conducts progressive model compression and fine-tuning. One of the most notable recent works is MiniLLM [465], which suggests using the reverse of the Kullback-Leibler Divergence to prevent student models from overestimating the low-probability distribution of the teacher.
@@
Black-box KD has recently garnered increased attention due to the rise of closed-source LLMs. The first category of black-box KD is ICL, initially introduced in GPT-3 [483]. In ICL, the teacher receives a task description, a few task examples, and a query, where the teacher predicts a response that the student aims to mimic. [484] combines in-context learning objectives with language modeling objectives to distill both the ability to interpret in-context examples and task knowledge into smaller models. AICD [486] explores the simultaneous distillation of in-context learning and reasoning, leveraging the autoregressive nature of LLMs to optimize the likelihood of all rationales in context. [485] introduces a novel framework to iteratively train dense retrievers capable of identifying high-quality in-context examples for LLMs.
@@
The second category of black-box KD is IF, which seeks to enhance the zero-shot capabilities of LLMs through instruction prompts. Specifically, in IF, the teacher generates task-specific instructions on which the student network is fine-tuned [481, 479, 482, 478, 480]. Self-Instruct [478] generates instructions, input, and output samples, filters out invalid samples, and then fine-tunes the student model. UniversalNer [479] explores mission-focused instruction tuning to train student models capable of excelling in a broad application class. Lion [481] utilizes an adversarial framework to generate challenging instructions for students, while LaMini-LM [480] compiles a large dataset containing 2.58 million instructions based on both existing and newly generated instructions.
@@
The last and most popular black-box KD approach is COT, where the teacher model generates rationales alongside predictions, providing the student network with intermediate inference steps. In COT, the student is expected to generate predictions and reasonings similar to the teacher, similar to feature learning where the student mimics the teacher’s intermediate steps for better predictions. This concept was first introduced by [466] and has been expanded upon in subsequent works [468, 470, 469, 467, 471, 472, 474, 473, 475, 476, 477]. [467] explores the trade-off between model and dataset size, demonstrating that the reasoning ability of LLMs can be transferred to SLMs. MCC-KD [469] generates multiple reasonings for each sample and enforces consistency and diversity among them. [470] employs a multi-task learning framework to train the student with the teacher’s reasonings. In contrast to other approaches that overlook reasonings with incorrect labels, [475] utilizes both positive and negative data. Most recently, [476] demonstrated that generating backward questions and reasoning, and training the student on both forward and backward reasoning, significantly enhances the generalization ability of the student.
@@
VI-B Foundation Model Distillation
@@
Prominent FMs, including VLMs such as CLIP [263], conversational agents such as ChatGPT [483, 489, 490], and generative models such as DALL-E [491], illustrate the diverse domains from which knowledge can be distilled using KD. KD is utilized to distill the complex, multi-modal representations from these models into simpler ones optimized for specific tasks. In addition, vision-specific models such as SAM [492] and DINO [493] serve as valuable sources of knowledge that can be distilled for various applications, including segmentation and unsupervised object discovery. Transformer-based models, including ViT [494], DETR [495], Swin Transformer [496], and DeIT [497] also provide foundational knowledge that can be transferred through KD to enhance performance in vision-related tasks.
@@
VI-B1 Vision-language Models and Knowledge Distillation
VLMs bridge visual and textual data through joint representation learning, offering capabilities to align these data types effectively. KD plays a central role in refining these models [498, 485, 499], enabling their general-purpose knowledge to be adapted for more compact or task-specific models, enhancing scalability and reducing computational overhead.
CLIP (Contrastive Language-Image Pre-training) [263], one of the most significant VLMs, exemplifies this integration. Trained on paired image and text data, CLIP achieves remarkable generalization by aligning global visual and textual representations. Through KD, these capabilities are transferred into student models designed for downstream tasks, including several key areas:
@@ VI-B1 Vision-language Models and Knowledge Distillation
In conventional vision tasks such as semantic segmentation and object detection, a dataset-focused methodology is traditionally employed, where the most successful techniques depend on a training dataset that has been manually annotated for a predefined and narrow range of categories. However, the rise of advanced VLMs, namely CLIP, is driving a shift towards an open-world paradigm. These models are trained through a simple yet scalable process that aligns image-text pairs, leveraging broad, loosely descriptive captions that can be collected in vast amounts with little human intervention. In object detection, KD techniques [500, 501, 502, 503, 504, 505, 506, 507, 508, 509, 510, 511, 512, 513, 514, 515, 516, 517, 506] distill CLIP’s global visual-textual knowledge into models capable of detecting objects outside fixed vocabularies, while in semantic segmentation [518, 519, 520, 521, 522, 523, 501, 524, 525, 526, 527, 528, 529, 530], and part segmentation [531] KD aids in transferring CLIP’s textual alignment to dense pixel-level tasks, enabling fine-grained scene understanding. Furthermore, [532] has explored open-vocabulary customization from CLIP through Data-Free Knowledge Distillation (DFKD), introducing a framework that enables CLIP model adaptation without requiring the original training data.
@@
The utilization of CLIP has recently been extensively explored for weakly supervised semantic segmentation [533, 534, 535, 536, 537, 538, 539, 540] where it enables the generation of segmentation masks with minimal labeled data, by leveraging its ability to align textual and visual representations within a shared embedding space. Through this approach, text prompts representing class labels guide the identification of relevant regions in an image by querying CLIP’s pre-trained image-text alignment capabilities. The visual features extracted by CLIP’s image encoder are matched to the semantic meaning encapsulated in text embeddings, producing coarse class activation maps that localize objects and regions associated with the given labels.
@@
Prompt learning is a method that enables the use of a large pre-trained model, such as CLIP, for specific tasks without having to retrain the entire model. This approach suggests adjusting the model’s representations for particular tasks by using learnable soft prompts, either text-based or visual, rather than relying on pre-designed hard prompts. Several methods [541, 542, 543, 544, 545] have been proposed to implement this approach.
@@
Applications such as generation and editing of 2D [546, 547, 548] and 3D images [549, 550, 551, 552, 553, 554, 555, 556], where text-guided modifications are required, also benefit significantly from CLIP-guided KD.
@@
The range of tasks addressed by KD in VLMs, particularly CLIP, is broad, including person re-identification [557], zero-shot human-object interaction (HOI) detection [558], open-vocabulary out-of-distribution classification [559], domain generalization [560], video-language retrieval [561], affordance detection [562], video highlight detection [563], and monocular depth estimation [564]. These applications demonstrate how KD transforms large-scale VLMs into accessible, efficient tools for specialized needs, with some works [565, 566, 567, 568] tackling multiple tasks simultaneously.
VI-B2 Segment Anything (SAM)
The Segment Anything Model (SAM) [492] is a versatile and general-purpose model designed for efficient and accurate object segmentation across diverse image domains. Developed with a focus on flexibility, SAM employs a powerful vision-transformer architecture that can segment any object in an image based on prompts, such as points, boxes, or masks. Its pre-trained nature enables SAM to generalize well across tasks without task-specific fine-tuning, making it a valuable resource for many applications. Some works [323, 569, 570] leverage SAM’s pre-trained knowledge by distilling its segmentation capabilities into lighter, task-specific models, enabling the transfer of its robust segmentation performance to applications with constrained data annotations, computational resources or domain-specific requirements.

### v2

@@ 5.4 Speech Input
Feature-based Wang et al.(Wang et al., 2024c), Liu et al.(Liu et al., 2019b), V2I-DETR (Jiang et al., 2024c), MaskAgain(Li et al., 2023e)
Similarity-based RSKD(Wang & Tang, 2023), Bhardwaj et al.(Bhardwaj et al., 2019), OOKD (Kim et al., 2024a), DL-DKD(Dong et al., 2023a)
Logit-based Camarena et al.(Camarena et al., 2024), Wu et al.(Wu & Chiu, 2020), PKD (Soufleri et al., 2024), MobileVOS(Miles et al., 2023), V2I-DETR(Jiang et al., 2024c), DTO(Monti et al., 2022), MVD (Wang et al., 2023l), RankDVQA-mini(Feng et al., 2024a), VideoAdviser (Wang et al., 2023o), Afouras et al.(Afouras et al., 2020), Perez et al.(Perez et al., 2020)
5.5 Video Input
Knowledge distillation has been widely applied in image and language tasks, but its potential in video-based applications is gaining increasing attention. Recent studies have explored KD for video action recognition (Liu et al., 2019b; Wu & Chiu, 2020; Camarena et al., 2024; Wang et al., 2024c; Soufleri et al., 2024), video classification (Bhardwaj et al., 2019; Wang & Tang, 2023), video object segmentation (Miles et al., 2023; Jiang et al., 2024c), video instance segmentation (Kim et al., 2024a), Partially Relevant Video Retrieval (PRVR) (Dong et al., 2023a), trajectory forecasting (Monti et al., 2022), and representation learning through masked video modeling (Li et al., 2023e; Wang et al., 2023l). Table 8 summarizes video distillation methods.
The main challenge in deep learning models, particularly in applications such as video quality estimation (Feng et al., 2024a) and attention prediction (Fu et al., 2020), is the high computational cost and complexity of large models, which restrict their execution on edge devices. KD effectively addresses this by enabling a smaller, more efficient model (student) to learn from a larger, complex model (teacher) with minimal performance reduction while significantly decreasing computational and memory requirements. MobileVOS (Miles et al., 2023) specifically focuses on semi-supervised video object segmentation on resource-constrained devices, for example mobile phones, and proposes a pixel-wise representation distillation loss to efficiently transfer structural information from the teacher to the student while ensuring feature consistency across frames. In Kim et al. (2024a), the authors propose a real-time approach that distills similarity information from an offline teacher model, which processes entire video sequences, to an online model, which processes individual frames. Mullapudi et al. (2019) introduces an online approach that applies feature distillation on live videos for low-cost semantic segmentation.
Furthermore, KD is applicable in multi-modal learning, enabling cross-modal knowledge transfer, such as video-to-image (Jiang et al., 2024c), video-to-text (Wang et al., 2023o), text-to-video (Dong et al., 2023a), audio-to-video (Afouras et al., 2020), and video-to-audio (Perez et al., 2020). Notably, (Pan et al., 2020) leverages KD in a spatio-temporal graph model for the video captioning task, making the model more robust against spurious correlations. V2I-DETR enhances medical video lesion detection by distilling temporal knowledge from a video-based teacher model to an image-based student model, achieving real-time inference speed with high accuracy (Jiang et al., 2024c).
6 Applications
@@
6.1 Distillation in Large Language Models
Distillation in Large Language ModelsBlack-boxWhite-boxChain of ThoughtInstruction FollowingIn-context LearningBERT-basedDecoder-based TinyBERT(Jiao et al., 2019), DistillBERT(Sanh, 2019), PKD(Sun et al., 2019), Turc et al.(Turc et al., 2019), MobileBERT(Sun et al., 2020), Minilm(Wang et al., 2020b), Xtremdistil(Mukherjee & Awadallah, 2020), Xtremedistiltransformer(Mukherjee et al., 2021), MetaDistill(Zhou et al., 2021b) TED(Liang et al., 2023a), GKD(Agarwal et al., 2023), Agarwal et al.(Agarwal et al., 2024), Pc-LoRA(Hwang et al., 2024), Minillm(Gu et al., 2024) Li et al.(Li et al., 2022e), Magister et al.(Magister et al., 2022), Fine-tune-CoT(Ho et al., 2022), MCC-KD (Chen et al., 2023a), Distilling step-by-step(Hsieh et al., 2023), CSoTD(Li et al., 2023d), SCOTT(Wang et al., 2023j), DOCTOR(Chae et al., 2023), PaD(Zhu et al., 2023b), TGD(Li et al., 2024g), RevThink(Chen et al., 2024c), MiniPLM(Zhuang et al., 2025) Self-instruct(Wang et al., 2022b), UniversalNER(Zhou et al., 2023c), LaMini-LM(Wu et al., 2023c), Lion(Jiang et al., 2023b), Li et al.(Li et al., 2024d) Brown et al.(Brown et al., 2020), ILD(Huang et al., 2022e), LLM-R(Wang et al., 2023h), AICD(Liu, 2024)
@@ 6.1 Distillation in Large Language Models
LLMs have recently revolutionized the field of NLP, with significant applications across various domains in both academia and industry. However, these models pose critical challenges due to their large number of parameters, which limits their deployment in real-time scenarios. Additionally, these giant models, with a considerable number of attention blocks, have been shown to be overparameterized (Kaplan et al., 2020; Fischer et al., 2024). Therefore, KD plays a crucial role in compressing LLMs into Small Language Models (SLMs). Existing KD methods for LLMs are mainly classified into two major categories: white-box distillation and black-box distillation. In white-box KD, logits and intermediate outputs of the LLMs are accessible, enabling the student to benefit from the internal information of the teacher. This information can be either logits (Hinton, 2015) or intermediate features (Romero et al., 2014), such as outputs of each attention block or attention scores. On the other hand, in black-box distillation, only the predictions of the teacher are available for distillation, which is common in many closed-source LLMs like GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), limiting the distillation process to only the predictions. Black-box methods are categorized into In-Context Learning (ICL) (Huang et al., 2022e), Chain of Thought (CoT) (Li et al., 2022e), and Instruction Following (IF) (Wang et al., 2022b). Figure 8 summarizes the black-box and white-box distillation methods for LLMs.
@@
Most white-box KD methods focus on the BERT (Kenton & Toutanova, 2019) model (Turc et al., 2019; Jiao et al., 2019; Sanh, 2019; Sun et al., 2019; 2020; Wang et al., 2020b; Mukherjee & Awadallah, 2020; Mukherjee et al., 2021). DistilBERT (Sanh, 2019) was one of the works that initialized the student with a pre-trained teacher and minimized the soft probabilities between the student and the teacher during the pre-training phase. In a concurrent work, MiniLM (Wang et al., 2020b) proposed distilling the self-attention module of the last transformer layer of the teacher. MobileBERT (Sun et al., 2020) first trains a special teacher and then uses feature and attention transfer to distill knowledge to the student. TinyBERT (Jiao et al., 2019) introduced a KD method for both pre-training and task-specific training of a smaller version of BERT by distilling the logits, attention matrices, hidden states, and the output of the teacher’s embedding layer. PKD (Sun et al., 2019) introduces two strategies for incremental distillation, while Mukherjee & Awadallah (2020) and Mukherjee et al. (2021) propose architecture-agnostic and task-agnostic distillation methods, respectively.
@@
Recently, several new white-box KD methods have been proposed (Zhou et al., 2021b; Agarwal et al., 2023; Liang et al., 2023a; Hwang et al., 2024; Agarwal et al., 2024; Gu et al., 2024; Anshumann et al., 2025), primarily designed for decoder-based LLMs. MetaDistill (Zhou et al., 2021b) utilizes a feedback mechanism within a meta-learning framework, GKD (Agarwal et al., 2023) trains the student on its self-generated output sequences, and PC-LoRA (Hwang et al., 2024) concurrently conducts progressive model compression and fine-tuning. One of the most notable recent works is MiniLLM (Gu et al., 2024), which suggests using the reverse of the Kullback-Leibler Divergence to prevent student models from overestimating the low-probability distribution of the teacher. CKA(Dasgupta & Cohn, 2025) matches hidden states of different dimensionality, allowing for smaller students and higher compression ratios, and Gong et al. (2025) Alignings feature dynamics for effective KD.
Furthermore, white-box KD has been successfully employed in industry to train small language models in pre-training stage (Liu et al., 2023b; Sreenivas et al., 2024; Muralidharan et al., 2024; Wang et al., 2024b; Team et al., 2025; Yang et al., 2025). Liu et al. (2023b) proposes a data-free, quantization-aware training approach in which a quantized version of the teacher is trained as a student using data-free distillation. Pruning-aware distillation has also been adopted by NVIDIA (Muralidharan et al., 2024; Sreenivas et al., 2024). Muralidharan et al. (2024) prunes the teacher’s neurons and uses a small portion of the original data for post-training the student via KD. In a related approach, Sreenivas et al. (2024) additionally introduces a teacher correction step prior to distillation. Gemma3 (Team et al., 2025) and Qwen3 (Yang et al., 2025) employ logit sampling and weak-to-strong distillation, respectively.
@@
Black-box KD has recently garnered increased attention due to the rise of closed-source LLMs. The first category of black-box KD is ICL, initially introduced in GPT-3 (Brown et al., 2020). In ICL, the teacher receives a task description, a few task examples, and a query, where the teacher predicts a response that the student aims to mimic. Huang et al. (2022e) combines in-context learning objectives with language modeling objectives to distill both the ability to interpret in-context examples and task knowledge into smaller models. AICD (Liu, 2024) explores the simultaneous distillation of in-context learning and reasoning, leveraging the autoregressive nature of LLMs to optimize the likelihood of all rationales in context. Wang et al. (2023h) introduces a novel framework to iteratively train dense retrievers capable of identifying high-quality in-context examples for LLMs.
@@
The second category of black-box KD is IF, which seeks to enhance the zero-shot capabilities of LLMs through instruction prompts. Specifically, in IF, the teacher generates task-specific instructions on which the student network is fine-tuned (Jiang et al., 2023b; Zhou et al., 2023c; Li et al., 2024d; Wang et al., 2022b; Wu et al., 2023c). Self-Instruct (Wang et al., 2022b) generates instructions, input, and output samples, filters out invalid samples, and then fine-tunes the student model. UniversalNer (Zhou et al., 2023c) explores mission-focused instruction tuning to train student models capable of excelling in a broad application class. Lion (Jiang et al., 2023b) utilizes an adversarial framework to generate challenging instructions for students, while LaMini-LM (Wu et al., 2023c) compiles a large dataset containing 2.58 million instructions based on both existing and newly generated instructions.
@@
The last and most popular black-box KD approach is COT, where the teacher model generates rationales alongside predictions, providing the student network with intermediate inference steps. In COT, the student is expected to generate predictions and reasonings similar to the teacher, similar to feature learning where the student mimics the teacher’s intermediate steps for better predictions. This concept was first introduced by Li et al. (2022e) and has been expanded upon in subsequent works (Ho et al., 2022; Hsieh et al., 2023; Chen et al., 2023a; Magister et al., 2022; Li et al., 2023d; Wang et al., 2023j; Zhu et al., 2023b; Chae et al., 2023; Li et al., 2024g; Chen et al., 2024c; Zhuang et al., 2025). Magister et al. (2022) explores the trade-off between model and dataset size, demonstrating that the reasoning ability of LLMs can be transferred to SLMs. MCC-KD (Chen et al., 2023a) generates multiple reasonings for each sample and enforces consistency and diversity among them. Hsieh et al. (2023) employs a multi-task learning framework to train the student with the teacher’s reasonings. In contrast to other approaches that overlook reasonings with incorrect labels, Li et al. (2024g) utilizes both positive and negative data. Most recently, Chen et al. (2024c) demonstrated that generating backward questions and reasoning, and training the student on both forward and backward reasoning, significantly enhances the generalization ability of the student.
@@
6.2 Foundation Model Distillation
@@
Prominent FMs, including VLMs such as CLIP (Radford et al., 2021), conversational agents such as ChatGPT (Brown et al., 2020; Radford, 2018; Baktash & Dawodi, 2023), and generative models such as DALL-E (Ramesh et al., 2021), illustrate the diverse domains from which knowledge can be distilled using KD. KD is utilized to distill the complex, multi-modal representations from these models into simpler ones optimized for specific tasks. In addition, vision-specific models such as SAM (Kirillov et al., 2023) and DINO (Caron et al., 2021) serve as valuable sources of knowledge that can be distilled for various applications, including segmentation and unsupervised object discovery. Transformer-based models, including ViT (Dosovitskiy, 2020), DETR (Carion et al., 2020), Swin Transformer (Liu et al., 2021e), and DeIT (Touvron et al., 2021) also provide foundational knowledge that can be transferred through KD to enhance performance in vision-related tasks.
@@
6.2.1 Vision-language Models and Knowledge Distillation
VLMs bridge visual and textual data through joint representation learning, offering capabilities to align these data types effectively. KD plays a central role in refining these models (Chen et al., 2024i; Wang et al., 2023h; Minderer et al., 2022), enabling their general-purpose knowledge to be adapted for more compact or task-specific models, enhancing scalability and reducing computational overhead.
CLIP (Contrastive Language-Image Pre-training) (Radford et al., 2021), one of the most significant VLMs, exemplifies this integration. Trained on paired image and text data, CLIP achieves remarkable generalization by aligning global visual and textual representations. Through KD, these capabilities are transferred into student models designed for downstream tasks, including several key areas:
@@ 6.2.1 Vision-language Models and Knowledge Distillation
In conventional vision tasks such as semantic segmentation and object detection, a dataset-focused methodology is traditionally employed, where the most successful techniques depend on a training dataset that has been manually annotated for a predefined and narrow range of categories. However, the rise of advanced VLMs, namely CLIP, is driving a shift towards an open-world paradigm. These models are trained through a simple yet scalable process that aligns image-text pairs, leveraging broad, loosely descriptive captions that can be collected in vast amounts with little human intervention. In object detection, KD techniques (Gu et al., 2021b; Ma et al., 2022; Bangalath et al., 2022; Zang et al., 2022; Xie & Zheng, 2022; Kim et al., 2023a; Wu et al., 2023d; Wang et al., 2023i; Du et al., 2022; Feng et al., 2022; Gao et al., 2022b; Huynh et al., 2022; Long et al., 2023; Zhao et al., 2022f; Lin et al., 2022a; Kuo et al., 2022; Zhou et al., 2022d; Yao et al., 2022) distill CLIP’s global visual-textual knowledge into models capable of detecting objects outside fixed vocabularies, while in semantic segmentation (Lüddecke & Ecker, 2022; Ding et al., 2022a; Li et al., 2022a; Zabari & Hoshen, 2021; Zhou et al., 2022a; 2023d; Ma et al., 2022; Liang et al., 2023b; Xu et al., 2022a; Ghiasi et al., 2022; Shin et al., 2022a; Qin et al., 2023; Wysoczańska et al., 2024a; Chen et al., 2023b), and part segmentation (Choi et al., 2024) KD aids in transferring CLIP’s textual alignment to dense pixel-level tasks, enabling fine-grained scene understanding. Furthermore, Wei et al. (2025) has explored open-vocabulary customization from CLIP through Data-Free Knowledge Distillation (DFKD), introducing a framework that enables CLIP model adaptation without requiring the original training data.
@@
The utilization of CLIP has recently been extensively explored for weakly supervised semantic segmentation (Xie et al., 2022; Lin et al., 2023b; Xu et al., 2023b; Deng et al., 2023; Murugesan et al., 2024; Zhang et al., 2024a; Zhu et al., 2024a; Jang et al., 2024; Zhou et al., 2025) where it enables the generation of segmentation masks with minimal labeled data, by leveraging its ability to align textual and visual representations within a shared embedding space. Through this approach, text prompts representing class labels guide the identification of relevant regions in an image by querying CLIP’s pre-trained image-text alignment capabilities. The visual features extracted by CLIP’s image encoder are matched to the semantic meaning encapsulated in text embeddings, producing coarse class activation maps that localize objects and regions associated with the given labels.
@@
Prompt learning is a method that enables the use of a large pre-trained model, such as CLIP, for specific tasks without having to retrain the entire model. This approach suggests adjusting the model’s representations for particular tasks by using learnable soft prompts, either text-based or visual, rather than relying on pre-designed hard prompts. Several methods (Zhou et al., 2022c; b; Li et al., 2024h; Ge et al., 2023; Rao et al., 2022) have been proposed to implement this approach.
@@
Applications such as generation and editing of 2D (Abdal et al., 2022; Gal et al., 2022; Patashnik et al., 2021) and 3D images (Hyung et al., 2023; Canfes et al., 2023; Hong et al., 2022a; Jain et al., 2022; Kobayashi et al., 2022; Michel et al., 2022; Sanghi et al., 2022; Wang et al., 2022a), where text-guided modifications are required, also benefit significantly from CLIP-guided KD.
@@
The range of tasks addressed by KD in VLMs, particularly CLIP, is broad, including person re-identification (Liu et al., 2024c), zero-shot human-object interaction (HOI) detection (Wan & Tuytelaars, 2024), open-vocabulary out-of-distribution classification (Li et al., 2023f), domain generalization (Huang et al., 2023c), video-language retrieval (Pei et al., 2023), affordance detection (Cuttano et al., 2024), video highlight detection (Han et al., 2024a), and monocular depth estimation (Son & Lee, 2024). These applications demonstrate how KD transforms large-scale VLMs into accessible, efficient tools for specialized needs, with some works (Yang et al., 2023a; Nair, 2024; Wu et al., 2023b; Dong et al., 2023b) tackling multiple tasks simultaneously.
6.2.2 Segment Anything (SAM)
The Segment Anything Model (SAM) (Kirillov et al., 2023) is a versatile and general-purpose model designed for efficient and accurate object segmentation across diverse image domains. Developed with a focus on flexibility, SAM employs a powerful vision-transformer architecture that can segment any object in an image based on prompts, such as points, boxes, or masks. Its pre-trained nature enables SAM to generalize well across tasks without task-specific fine-tuning, making it a valuable resource for many applications. Some works (Zhang et al., 2024e; Hetang et al., 2024; Rahman et al., 2024) leverage SAM’s pre-trained knowledge by distilling its segmentation capabilities into lighter, task-specific models, enabling the transfer of its robust segmentation performance to applications with constrained data annotations, computational resources or domain-specific requirements.